OpenStudy (trojanpoem):

Find the limit (No L'hopital rule): written down. limx→0sinx1−2sin(x+π6)

OpenStudy (trojanpoem):

0/0

OpenStudy (trojanpoem):

Bro, I tried all this.

OpenStudy (irishboy123):

lol!! i will have a go too, interesting

OpenStudy (trojanpoem):

sin(x+30) = sinxcos30 + sin30cosx 1-2sin(x+30) = 1- sqrt(3)sinx + cosx sinx/ 1- sqrt(3)sinx + cosx

OpenStudy (trojanpoem):

Give it a go, and tell me what was the result.

OpenStudy (rational):

as it stands, im finding it difficult to interpet your function can you use latex and put it in more friendly form ?

OpenStudy (trojanpoem):

\[\lim_{x \rightarrow 0} \frac{ sinx }{ 1-2\sin(x+\frac{ \pi }{ 6 }) }\]

OpenStudy (irishboy123):

break it up as you did before (but correct the sin on the bottom) ie it should be sinx/ 1- sqrt(3)sinx - cosx and limit should be -1/ root(3)

OpenStudy (trojanpoem):

right it's -1/sqrt(3) what did you do ?

OpenStudy (irishboy123):

exactly what you did. break it up and work through. you get to x/-root(3)x, right? lim x/x = 1

OpenStudy (trojanpoem):

I can't.

OpenStudy (trojanpoem):

@IrishBoy123

OpenStudy (irishboy123):

\[\frac{sinx}{1- 2\sin(x) \cos \frac{π}{6} - 2\cos(x) \sin \frac{π}{6}}\] \[\frac{sinx}{1 - 2 \sin(x) \frac {\sqrt 3}{2} - 2 \cos(x) \frac{1}{2}}\] \[\frac{x}{1 - \sqrt{3} x - 1}\]

OpenStudy (trojanpoem):

sinx1−2sin(x)3√2−2cos(x)12 sinx/ 1- sqrt(3)sinx - cosx last step ?

OpenStudy (trojanpoem):

@IrishBoy123 , What kind of magic did you do to get x/1-sqrt3)x - 1

OpenStudy (irishboy123):

sin x ->x, cos x ->1 as x ->0

OpenStudy (trojanpoem):

Right the steps.

OpenStudy (irishboy123):

those are the steps, there is little else too it apart maybe from finding the limit symbol in the equation drawing thingy, which i have never been able to do and stuffing that on the LHS of the last one before introducing the limits. here are some more: http://planetmath.org/listofcommonlimits

OpenStudy (irishboy123):

actually let me try \[\lim_{x \rightarrow 0} \frac{sinx}{x} = 1\] et voila!

OpenStudy (trojanpoem):

sinx/x / 1/x - sinx/x * sqrt(3) - cosx/x now ??? 1/ 1/0 - 1/0 undefined - undefined

OpenStudy (irishboy123):

OK \[\lim_{x \rightarrow 0} \frac{sinx}{1 - \sqrt{3}sinx - \cos x} = \frac{x}{1 - \sqrt{3}x - 1} = -\frac{x}{\sqrt{3}x} = - \frac{1}{\sqrt{3}}\]

OpenStudy (trojanpoem):

Bro, as soon as you apply that sinx/x =1 you must put each x =0. You can't multiply by x.

OpenStudy (irishboy123):

these are limits!

OpenStudy (trojanpoem):

The final result is right, but the steps are defected.

OpenStudy (phi):

the limit of a quotient is the quotient of the limits in other words, you can do this \[ \lim_{x \rightarrow 0} \frac{\sin x}{1 - \sqrt{3}\sin x - \cos x} \\ \frac{ \lim_{x \rightarrow 0} \sin x} {\lim_{x \rightarrow 0}( 1 - \sqrt{3}\sin x - \cos x)} \] also, the limit of a sum is the sum of the limits

OpenStudy (irishboy123):

@TrojanPoem "The final result is right, but the steps are defected. " you are the re-incarnation of Bishop Berkeley?!?! right ?!?!

OpenStudy (trojanpoem):

@IrishBoy123 , Excuse me, thanks. @phi, Thanks , I though I can't do that.

OpenStudy (phi):

When you have time, see http://ocw.mit.edu/resources/res-18-006-calculus-revisited-single-variable-calculus-fall-2010/part-i-sets-functions-and-limits/lecture-5-a-more-rigorous-approach-to-limits/ (you might want to also watch the lecture previous to this one, which provides the background and lots of intuition)

OpenStudy (trojanpoem):

Thanks that's what I needed.

OpenStudy (phi):

to figure out the limit without using L'Hopital, we could use series expansions once you get to \[ \lim_{x \rightarrow 0} \frac{\sin x}{1 - \sqrt{3}\sin x - \cos x} \] I would rewrite it as \[ \lim_{x \rightarrow 0} \frac{\frac{\sin x}{x}}{\frac{1 - \cos x}{x} - \sqrt{3}\frac{\sin x}{x}} \] and now take the limit of the top and bottom the limit of the numerator is \[ \lim_{x \rightarrow 0} \frac{\sin x}{x} = 1\] the limit of the bottom is \[ \lim_{x \rightarrow 0}\frac{1 - \cos x}{x} -\sqrt{3}\lim_{x \rightarrow 0}\frac{\sin x}{x} \] as before the second term has a limit of -sqr(3)*1 for the first term, use \[ \cos(x) = 1 - \frac{x^2}{2!} + \frac{x^4}{4!} - ...\] so \[ 1 - \cos x= \frac{x^2}{2!} - \frac{x^4}{4!}+ \frac{x^6}{6!} - ...\] and \[ \frac{1- \cos x}{x} = \frac{x}{2!} - \frac{x^3}{4!}+ \frac{x^5}{6!} - ...\] now take the limit \[ \lim_{x \rightarrow 0}\frac{1 - \cos x}{x} =\lim_{x \rightarrow 0} \left( \frac{x}{2!} - \frac{x^3}{4!}+ \frac{x^5}{6!} -...\right) =0 \] putting it all together \[ \lim_{x \rightarrow 0} \frac{\frac{\sin x}{x}}{\frac{1 - \cos x}{x} - \sqrt{3}\frac{\sin x}{x}} = \frac{1}{0 - \sqrt{3} \cdot 1}= - \frac{\sqrt{3}}{3}\]

OpenStudy (phi):

another way to find the limit of \[ \lim_{x \rightarrow 0}\frac{1 - \cos x}{x} \\ = \lim_{x \rightarrow 0}\frac{1 - \cos x}{x} \frac{(1+\cos x)}{(1+\cos x)} \\ = \lim_{x \rightarrow 0}\frac{1 - \cos^2 x}{x(1+\cos x)} \\ = \lim_{x \rightarrow 0}\frac{\sin x}{x} \frac{\sin x}{(1+\cos x)} \] and now we can do \[ \lim_{x \rightarrow 0}\frac{\sin x}{x} \lim_{x \rightarrow 0}\frac{\sin x}{(1+\cos x)} \] to get 1 * 0 = 0

OpenStudy (trojanpoem):

Thanks, I didn't take the first one which involve 4!